Describe the transformation of the graph f(x) = 6x that yields the graph of (x) = 6x+1
Answer #1

46)=63 Let f(x) be y y=6* x yor f(x F(x)=62 goaph ATX -3 -2 V 2 3 1 1 -13 for f(x)=6+1 everything shifts by one unit X -117 Те > 9: 6xt ) ? > 926x 9ар, 1 2 3 , Т-6 세 1-13 It is slightly shifted upward by one unit.